An alpha particle with kinetic energy 11.0 MeV makes a collision with lead nucleus, but it is not "aimed" at the center of the lead nucleus, and has an initial nonzero angular momentum (with respect to the stationary lead nucleus) of magnitude L=p0b, where p0 is the magnitude of the initial momentum of the alpha particle and b=1.50×10−12 m . (Assume that the lead nucleus remains stationary and that it may be treated as a point charge. The atomic number of lead is 82. The alpha particle is a helium nucleus, with atomic number 2.)

a) What is the distance of closest approach?

r =

1.51×10−12

m

b) Repeat for b=1.30×10−13 m .

c)Repeat for b=1.20×10−14 m .
Answer #1

Here

mv1b = mv2r

where v2 = v1 * (b / r)

Therefore, using conservation of energy, we have

E = 1/2mv22 + kq1q2 / r

putting in the value of v2, we get

E = 1/2 * m * v12 * (b2 / r2) + kq1q2 / r

solving this, we get a quadratic equation

Er2 - kq1q2*r - Eb2 = 0

where we can find r.

so,

Let's put b = 1.50e-12 m

E = 11 MeV = 11e6 * 1.6e-19 = 1.76e-12 J

q1 = 2e (Helium - alpha)

q2 = 82e (lead)

so, we have

1.76e-12 * r2 - 3.778e-26 * r - 3.96e-36 = 0

r = 3.778e-26 +/- sqrt ( 2.787e-47) / 2*1.76e-12

r = 3.778e-26 +/- 5.28e-24 / 3.52e-12

r = 1.51e-12 m

------------------------------------------------------------------------------------------

Now , for b = 1.30e-13 m

exactly same procedure

1.76e-12 * r2 - 3.778e-26 * r - 2.97e-38 = 0

r = 3.778e-26 +/- sqrt ( 2.105e-49) / 2*1.76e-12

r = 1.41e-13 m

-----------------------------------------------------------------------------------------

For b = 1.20×10−14

1.76e-12 * r2 - 3.778e-26 * r - 2.53e-40 = 0

r = 3.778e-26 +/- sqrt ( 3.208e-51) / 2*1.76e-12

r = 2.68e-14 m